# Differentially Private Adaptation of Diffusion Models via Noisy Aggregated Embeddings

### v1

@@ 2.3 Differentially Private Adaptation of Diffusion Models
Recent advancements in applying differential privacy (DP) to diffusion models have aimed to balance privacy preservation with the high utility of generative outputs. Dockhorn et al. dockhorn2023differentially proposed a differentially private diffusion model (DP diffusion) that enables privacy-preserving generation of realistic samples, setting a foundational approach for adapting diffusion processes under DP constraints. Another common strategy involves training a model on a large public dataset, followed by differentially private finetuning on a private dataset, as explored by Ghalebikesabi et al. ghalebikesabi2023differentially . While effective in certain contexts, this approach raises privacy concerns, particularly around risks of information leakage during the finetuning phase pmlr-v235-tramer24a .
In response to these limitations, various adaptation techniques have emerged. Although not specific to diffusion models, some methods focus on training models on synthetic data followed by DP-constrained finetuning, as in the VIP approach yu2024vip , which demonstrates the feasibility of applying DP in later adaptation stages. Other approaches explore differentially private learning of feature representations sander2024differentially , aiming to distill private information into a generalized embedding space while maintaining DP guarantees. Although these adaptations are not yet implemented for diffusion models, they lay essential groundwork for developing secure and efficient privacy-preserving generative models.
3 Differentially Private Adaptation via Noisy Aggregated Embeddings
3.1 General Approach
Let $x^{(1)},\dots,x^{(n)}$ represent a target dataset of images whose characteristics we wish to privately adapt our image generation towards. Assuming that we can encode each image $x^{(i)}$ as a directional embedding vector $u^{(i)}$, we can aggregate the embeddings $u^{(1)},\dots,u^{(n)}$ by calculating the centroid. The purpose of this aggregation is to limit the sensitivity of the final output to each $x^{(i)}$. In order to provide DP guanrantees, we also add isotropic Gaussian noise to the centroid. We can therefore define the resulting embedding vector $u^{*}$ as follows:
$u^{*}=\frac{1}{n}\sum_{i=1}^{n}u^{(i)}+\mathcal{N}(0,\sigma^{2}I)$ (7)
where the minimum $\sigma$ required to provide $(\varepsilon,\delta)$-DP is given by the following expression based on (balle2018improving, , Theorem 1):
$\sigma=\frac{\Delta}{n}\cdot\frac{\sqrt{2\log(1.25)/\delta}}{\varepsilon}.$ (8)
@@ 3.1 General Approach
The noisy centroid embedding $u^{*}$ can then be used to adapt the downstream image generation process. In Sections 3.3 and 3.4, we describe how to apply our method to TI and SG respectively, outlining the encoding of $x^{(i)}$ as $u^{(i)}$ and the role of $u^{*}$ in the image generation process. To reduce the amount of noise needed to provide the same level of DP, we employ subsampling: instead of computing the centroid over all $n$ embedding vectors, we randomly sample $m\leq n$ embedding vectors without replacement and compute the centroid over only the sampled vectors. Then the standard privacy amplification by subsampling bounds (such as (6)) can be applied. Formally, we sample $D_{\text{sub}}\subseteq\{u^{(1)},\dots,u^{(n)}\}$ where $|D_{\text{sub}}|=m$, and compute the output embedding as follows:
@@
3.2 Using Style Guidance
Applying our method to style guidance is straightforward. To obtain the embeddings, we simply use the CLIP image encoder to obtain the embedding for each image. More formally, for each $x^{(i)}$, let $u^{(i)}=\textsf{CLIP}(x^{(i)})$. We can then aggregate the embeddings $u^{(i)}$ to obtain $u^{*}$ via Equation (7) or (9), depending on whether we wish to incorporate subsampling. Finally, we use $u^{*}$ to guide the reverse diffusion process, as outlined in Section 2.1.1, by computing $\hat{\epsilon}_{\theta}(x_{t},y,t)$ via a modification of Equation (2):
$\hat{\epsilon}_{\theta}(x_{t},y,t)=\epsilon_{\theta}(x_{t},y,t)+w\sqrt{1-\alpha_{t}}\nabla_{x_{t}}\ell_{\cos}(u^{*},\textsf{CLIP}(\hat{x}_{0}))$ (10)
3.3 Using Textual Inversion
Textual Inversion (TI) is inherently parameter-efficient and offers certain privacy benefits, as information from an entire dataset of images is compressed into a single token embedding vector. This compression limits the model’s capacity to memorize specific images, making data extraction attacks difficult. However, this privacy is merely heuristic and yet to be proven, so TI may still be vulnerable to privacy attacks such as membership inference. A similar adapation technique with privacy guarantees may therefore be desirable.
To incorporate our approach into TI, instead of training a single token embedding on the entire dataset, we instead train a separate embedding $u^{(i)}$ on each $x^{(i)}$ to obtain a set of embeddings $u^{(1)},\dots,u^{(n)}$. We can formalize the encoding process as follows:
$u^{(i)}=\arg\min_{u}\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I),t}[\|\epsilon-\epsilon_{\theta}(x_{t}^{(i)},y_{u},t)\|^{2}]$ (11)
We can then aggregate the embeddings $u^{(i)}$ to obtain $u^{*}$ via Equation (7) or (9), depending on whether we wish to utilize subsampling. Similar to TI’s $u_{*}$, we can use $u^{*}$ to represent a new placeholder token $S^{*}$ that can be incorporated into prompts for personalized image generation. While $u^{*}$ may not fully solve the TI optimization problem presented in Equation (5), it provides provable privacy guarantees, with only a minimal trade-off in accurately representing the content or style of the training data.
@@ 4.1 Datasets
We compiled two datasets for our style adaptation experiments, designed to proxy for private datasets unrecognized by the base Stable Diffusion model. The first dataset consists of 158 artworks created by the artist @eveismyname, who has provided consent for non-commercial use. This dataset serves to examine how models can capture and replicate artistic styles without directly copying individual works. Although we acknowledge the possibility that a few of these artworks may have been inadvertently included in the pre-training of Stable Diffusion due their public presence on social media platforms, the artist’s limited recognition and modest portfolio size make it unlikely that their unique style could be effectively extracted from the pre-trained model. The second dataset consists of 47 pictograms from the Paris 2024 Olympics Paris2024Pictograms , approved strictly for non-commercial editorial use IOCProperties . These pictograms were unveiled in February 2023, several months after the release of Stable Diffusion v1.5, which serves as the base model for our experiments. Through this dataset, we hope to evaluate how effectively our approach adapts to novel and unfamiliar visual patterns. Representative samples from both datasets are presented in Figure 1.
4.2 Style Guidance Results
We applied our SG-based bansal2023universal approach, detailed in Section 3.2 to both datasets. While this method successfully provided privacy protection by obfuscating embedding details, the resulting image quality and style capture was notably less compared to the TI-based approach. The CLIP embeddings, even without noise or subsampling, led to images that retained only generalized stylistic elements, lacking the detailed fidelity and coherence achieved with TI. Results of our attempt can be seen in Figure 2. This result underscores the superiority of our TI method for maintaining both privacy and high-quality image generation.
A possible explanation for the difference in style transfer effectiveness is that SG, as with other guidance methods, suffers from instability and sensitive dependency to hyperparameters such as guidance weight $w$. While Bansal et al. proposed using backward guidance and per-step self-recurrence to remedy these issues, we found that they were insufficient for our application. Another possible explanation is that the stylistic signal of the image embedding vectors $\textsf{CLIP}(x^{(i)})$ are simply not resilient to our aggregation process.
@@ 4.2 Style Guidance Results
4.3 Textual Inversion Results
Figure 4: Images generated by Stable Diffusion v1.5 using the prompt “Icon of the Eiffel Tower in the style of <Paris 2024 Pictograms>”, with the embedding <Paris 2024 Pictograms> trained using different values of $m$ and $\varepsilon$.
Using both the @eveismyname and Paris 2024 pictograms dataset, we trained TI gal2023an embeddings on Stable Diffusion v1.5 rombach2022high using our proposed method detailed in Section 3.3. Our primary goal is to investigate how DP configurations, specifically the $\varepsilon$ and subsampling sizes ($m$), affect the generated images’ quality and privacy resilience. For regular TI, we utilize Stable Diffusion’s default process to embed the private dataset without any additional noise. For the DP-enhanced experiments using our method, we test multiple configurations of $m$ and $\varepsilon$ to analyze the trade-off between image fidelity and privacy resilience.
Figures 3 and  4 present generated images across three key configurations: (1) regular TI without DP, (2) centroid TI with DP at various epsilon values, and (3) centroid TI with DP at different subsampling sizes within the private dataset. We used the same random seed to generate embeddings and sample the images for ease of visual comparision between different configurations. As with common practice, we set $\delta=1/n$. Since $\sigma$ is undefined for $\varepsilon=0$, we demonstrate the results of $\varepsilon\approx 0$, in other words, infinite noise, by setting $\varepsilon=10^{-5}$. Images generated without DP closely resemble the unique stylistic elements of the target dataset. In particular, images adapted using @eveismyname images displayed crisp details and nuanced color gradients characteristic of the artist’s work, while those of Paris 2024 pictograms captured the logo’s original structure. In contrast, DP configurations introduce a discernible degradation in image quality, with lower epsilon values and smaller subsampling sizes resulting in more noticeable noise and diminished stylistic fidelity.
As $\varepsilon\rightarrow 0$, the resulting token embedding $u^{*}$ gradually loses its semantic meaning, leading to a loss of stylistic fidelity. In particular, $y_{u^{*}}$ tends towards $y$ (a conditioning vector independent of the learnable embedding). In our results, this manifests as a painting of Taylor Swift devoid of the artist-specific stylistic elements, or a generic icon of the Eiffel Tower (with color, as opposed to the black and white design of the pictograms). With this in mind, $\varepsilon$ can be interpreted as a drift parameter, representing the progression from the optimal $u^{*}$ towards infinity, gradually steering the generated image away from the target style in exchange for stronger privacy guarantees.
Meanwhile, reducing $m$ also reduces the sensitivity of the generated image to $\varepsilon$, as evident by the observation that, on both datasets at $m=4$, (subsampling rate below $0.1$) image generation can tolerate $\varepsilon$ as low as $0.5$ without significant changes in visual characteristics, and retaining stylistic elements of the target dataset at $\varepsilon$ as low as $0.1$. This strong boost in robustness comes at a small price of base style capture fidelity. As observed in figures 3 and 4, we can also treat subsampling as an introduction of noise. Mathematically, the subsample centroid is an unbiased estimate of the true centroid, and so the subsampling process itself defines a distribution centered at the true centroid. However, the amount of noise introduced by the subsampling process is limited by the individual image embeddings, as a subsample centroid can only stray from the true centroid as much as the most anomalous point in the dataset.

### v2

@@ 2.3 Differentially Private Adaptation of Diffusion Models
Recent advancements in applying DP to diffusion models have aimed to balance privacy preservation with the high utility of generative outputs. Dockhorn et al. (Dockhorn et al., 2023) proposed a Differentially Private Diffusion Model (DPDM) that enables privacy-preserving generation of realistic samples, setting a foundational approach for adapting diffusion processes using DP-SGD. Another common strategy involves training a model on a large public dataset, followed by differentially private fine-tuning on a private dataset, as explored by Ghalebikesabi et al. (2023). While effective in certain contexts, this approach raises privacy concerns, particularly around risks of information leakage during the fine-tuning phase (Tramèr et al., 2024).
In response to these limitations, various adaptation techniques have emerged. Although not specific to diffusion models, some methods focus on training models on synthetic data followed by DP-constrained fine-tuning, as in the VIP approach (Yu et al., 2024), which demonstrates the feasibility of applying DP in later adaptation stages. Other approaches explore differentially private learning of feature representations (Sander et al., 2024), aiming to distill private information into a generalized embedding space while maintaining DP guarantees. Although these adaptations are not yet implemented for diffusion models, they lay essential groundwork for developing secure and efficient privacy-preserving generative models.
3 Differentially Private Adaptation via Textual Inversion
TI is inherently parameter-efficient and offers certain privacy benefits, as information from an entire dataset of images is compressed into a single token embedding vector. This compression limits the model’s capacity to memorize specific images, making data extraction attacks difficult. However, this privacy is merely heuristic and yet to be proven, so TI may still be vulnerable to privacy attacks such as membership inference. A similar adaptation technique with privacy guarantees may therefore be desirable.
Let $x^{(1)},\dots,x^{(n)}$ represent a target dataset of images whose characteristics we wish to privately adapt our image generation towards. Instead of training a single token embedding on the entire dataset as in regular TI, we train a separate embedding $u^{(i)}$ on each $x^{(i)}$ to obtain a set of embeddings $u^{(1)},\dots,u^{(n)}$, as illustrated in Figure 1. We can formalize the encoding process as follows:
$u^{(i)}=\arg\min_{u}\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I),t}[\|\epsilon-\epsilon_{\theta}(x_{t}^{(i)},y_{u},t)\|^{2}].$ (4)
Then, we can aggregate the embeddings $u^{(1)},\dots,u^{(n)}$ by calculating the centroid. The purpose of this aggregation is to limit the sensitivity of the final output to each $x^{(i)}$. In order to provide DP guanrantees, we also add isotropic Gaussian noise to the centroid. We can therefore define the resulting embedding vector $u^{*}$ as follows:
$u^{*}=\frac{1}{n}\sum_{i=1}^{n}u^{(i)}+\mathcal{N}(0,\sigma^{2}I),$ (5)
where the minimum $\sigma$ required to provide $(\varepsilon,\delta)$-DP is given by the following expression based on Balle & Wang (2018, Theorem 1):
$\sigma=\frac{\Delta}{n}\cdot\frac{\sqrt{2\ln(1.25/\delta)}}{\varepsilon}.$ (6)
@@ 3 Differentially Private Adaptation via Textual Inversion
The noisy centroid embedding $u^{*}$ can then be used to adapt the downstream image generation process. Similar to regular TI’s $u_{*}$, we can use $u^{*}$ to represent a new placeholder token $S^{*}$ that can be incorporated into prompts for personalized image generation. While $u^{*}$ may not fully solve the TI optimization problem presented in (2), it provides provable privacy guarantees, with only a minimal trade-off in accurately representing the style of the target dataset.
To reduce the amount of noise needed to provide the same level of DP, we employ subsampling: instead of computing the centroid over all $n$ embedding vectors, we randomly sample $m\leq n$ embedding vectors without replacement and compute the centroid over only the sampled vectors. Then the standard privacy amplification by subsampling bounds (such as (3)) can be applied. Formally, we sample $D_{\text{sub}}\subseteq\{u^{(1)},\dots,u^{(n)}\}$ where $|D_{\text{sub}}|=m$, and compute the output embedding as follows:
@@ 4.1 Datasets
We compiled two datasets to evaluate our style adaptation method, specifically selecting content unlikely to be recognized by Stable Diffusion v1.5, our base model.
The first dataset consists of 158 artworks by the artist @eveismyname, who has granted consent for non-commercial use. This dataset allows us to assess whether models can capture artistic styles without memorizing individual works. While some of these artworks may have been publicly accessible on social media, making incidental inclusion in Stable Diffusion’s pretraining possible, the artist’s limited recognition and relatively small portfolio reduce the likelihood that the model has internalized her unique style. This dataset serves as a controlled test for privacy-preserving style transfer on individual artistic collections.
The second dataset contains 47 pictograms from the Paris 2024 Olympics (Paris 2024, ), permitted strictly for non-commercial editorial use (International Olympic Committee, ). These pictograms were officially released in February 2023, several months after the release of Stable Diffusion v1.5, ensuring they were absent from the model’s pretraining data. This dataset allows us to assess how well our approach adapts to newly introduced visual styles that the base model has never encountered.
Both datasets are used to test the ability of our method to extract and transfer stylistic elements while preserving privacy. Representative samples are shown in Figure 2.
4.2 Style Transfer Results
Figure 4: Images generated by Stable Diffusion v1.5 using the prompt “Icon of a dragon in the style of $<$Paris 2024 Pictograms$>$”, with the embedding $<$Paris 2024 Pictograms$>$ trained using different values of $m$ and $\varepsilon$.
Using both the @eveismyname and Paris 2024 pictograms dataset, we trained TI (Gal et al., 2023) embeddings on Stable Diffusion v1.5 (Rombach et al., 2022) using DPAgg-TI. Our primary goal is to investigate how DP configurations, specifically the privacy budget $\varepsilon$ and subsampling size $m$, affect the generated images quality and privacy resilience. For regular TI, we utilize the default process to embed the private dataset without any additional noise. For the DPAgg-TI, we test multiple configurations of $m$ and $\varepsilon$ to analyze the trade-off between image fidelity and privacy.
Figures 3 and  4 present generated images across two key configurations: (1) regular TI without DP, (2) DPAgg-TI with DP at different values of $m$ and $\varepsilon$. We used the same random seed to generate embeddings, subsample images, and sample DP noise for ease of visual comparision between different configurations. As with common practice, we set $\delta=1/n$. Since $\sigma$ is undefined for $\varepsilon=0$, we demonstrate the results of $\varepsilon\approx 0$, in other words, infinite noise, by setting $\varepsilon=10^{-5}$. The purpose of this parameter value is to demonstrate the image generated when $u^{*}$ contains zero information about the target dataset. Images generated without DP closely resemble the unique stylistic elements of the target dataset. In particular, images adapted using @eveismyname images displayed crisp details and nuanced color gradients characteristic of the artist’s work, while those of Paris 2024 pictograms captured the logo’s original structure. In contrast, DP configurations introduce a discernible degradation in image quality, with lower epsilon values and smaller subsampling sizes resulting in more noticeable noise and diminished stylistic fidelity.
As $\varepsilon\rightarrow 0$, the resulting token embedding $u^{*}$ gradually loses its semantic meaning, leading to a loss of stylistic fidelity. In particular, $y_{u^{*}}$ tends towards $y$ (a conditioning vector independent of the learnable embedding). In our results, this manifests as a painting of Taylor Swift devoid of the artist-specific stylistic elements, or a generic icon of a dragon (with color, as opposed to the black and white design of the pictograms). With this in mind, $\varepsilon$ can be interpreted as a drift parameter, representing the progression from the optimal $u^{*}$ towards infinity, gradually steering the generated image away from the target style in exchange for stronger privacy guarantees. We also observe instances where there is a temporary drop in prompt fidelity (e.g., $m=16,\varepsilon\in[0.5,1]$ in Figure 3 and intermediate $\varepsilon$ values in Figure 4) which restores as $u^{*}$ drifts even further from its optimal value. We hypothesize that this is due to drifted $u^{*}$ capturing a different meaning unrelated to the prompt, before losing any meaning that could be interpreted by Stable Diffusion’s text encoder, causing $u^{*}$ to be disregarded from $y_{u^{*}}$ and the prompt fidelity to be restored. Another possible explanation is that the temporary drop in prompt fidelity is due to the drift path of $u^{*}$ passing through non-linear regions within embedding space. We leave further investigations into this observation for future work.
Meanwhile, reducing $m$ also reduces the sensitivity of the generated image to $\varepsilon$, as evident by the observation that, on both datasets at $m=4$, (subsampling rate below $0.1$) image generation can tolerate $\varepsilon$ as low as $0.5$ without significant changes in visual characteristics, and retaining stylistic elements of the target dataset at $\varepsilon$ as low as $0.1$. This strong boost in robustness comes at a small price of base style capture fidelity. As observed in Figures 3 and 4, we can also treat subsampling as an introduction of noise. Mathematically, the subsample centroid is an unbiased estimate of the true centroid, and so the subsampling process itself defines a distribution centered at the true centroid. However, the amount of noise introduced by the subsampling process is limited by the individual image embeddings, as a subsample centroid can only stray from the true centroid as much as the biggest outlier in the dataset.
4.3 Quantitative Evaluation
4.3.1 User Study
To evaluate the utility of our approach under different DP and subsampling configurations, we conducted a user study with 25 participants. Each participant was shown reference images from the target dataset and asked to compare pairs of generated images, selecting the one that better captured the style of the reference images. Images were generated using 10 prompts and adapted TI embeddings for the @eveismyname and Paris 2024 Pictogram datasets, resulting in 20 groups of images. Each participant evaluated two groups, one randomly selected from each dataset, with comparisons focusing on model configurations differing by DP noise and subsampling size.
Survey results, summarized in Table 3 in Appendix A, align with our design goals. Participants showed no clear preference between regular TI and DPAgg-TI, suggesting that our privacy-preserving approach maintains perceptual quality. As expected, both DP noise and reduced subsampling size degraded style fidelity, consistent with the trade-offs inherent in differential privacy. Preferences at $\varepsilon=1$ were split, but subsampling was generally favored, reinforcing its role in reducing noise impact while preserving style.
4.3.2 Kernel Inception Distance
The Kernel Inception Distance (KID) (Bińkowski et al., 2018) is a metric for evaluating generative models by measuring the difference between the distributions of generated and training images in an embedding space. To compute KID, images generated by the model and real training images are passed through an Inception network (Szegedy et al., 2015), and their distributional differences are estimated. Unlike the more commonly used Fréchet Inception Distance (FID) (Heusel et al., 2017), KID is an unbiased estimator of the true divergence between the learned and target distributions (Jayasumana et al., 2024), making it more suitable for smaller datasets, as in our case.
We report KID scores for different parameters in Tables 2 and 2, showing that DPAgg-TI maintains the style transfer fidelity of TI while ensuring differential privacy. Further discussion of these results is provided in Appendix B.
Table 1: KID scores of DPAgg-TI on @eveismyname dataset for various $\varepsilon$ values ranging from $\varepsilon=10^{-5},0.1,0.5,1.0,2.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl).
$m$ No DP $\varepsilon=$ 5.0 $\varepsilon=$ 2.0 $\varepsilon=$ 1.0 $\varepsilon=$ 0.5 $\varepsilon=$ 0.1 $\varepsilon\approx$ 0
0.0444 0.0794 0.0422 0.0529 0.0690 0.1117 0.0654
32 0.0752 0.0845 0.0865 0.1167 0.0300 0.0649 0.0657
16 0.0351 0.0379 0.0430 0.0663 0.1309 0.0438 0.0658
8 0.0359 0.0366 0.0350 0.0366 0.0396 0.0530 0.0658
4 0.0245 0.0250 0.0249 0.0250 0.0258 0.0314 0.0653
ctrl 0.0318
Table 2: KID scores of DPAgg-TI on Paris dataset for various $\varepsilon$ values ranging from $\varepsilon=1e-5,0.1,0.5,1.0,2.0,5.0$ (including no DP) under different subsampling levels ($m=4,8,16,32$) as well as regular TI (ctrl).
$m$ No DP $\varepsilon=$ 5.0 $\varepsilon=$ 2.0 $\varepsilon=$ 1.0 $\varepsilon=$ 0.5 $\varepsilon=$ 0.1 $\varepsilon\approx$ 0
0.1146 0.1202 0.1368 0.1314 0.1389 0.1209 0.1274
32 0.1220 0.1036 0.1258 0.1377 0.1307 0.1245 0.1259
16 0.1311 0.1424 0.1170 0.1311 0.1381 0.1335 0.1278
8 0.1317 0.1307 0.1220 0.1117 0.1295 0.1313 0.1272
4 0.1141 0.1094 0.1137 0.1190 0.1194 0.1583 0.1259
ctrl 0.1388
4.4 Ablation Study
4.4.1 Textual Inversion with DP-SGD
Figure 5: Comparing our approach to applying DP-SGD to regular TI using prompt “a painting of Taylor Swift in the style of @eveismyname”.
A natural question that arises is how well our approach compares to the naive method of applying DP-SGD to regular TI training. We therefore integrated DP-SGD into the TI codebase using the Opacus library and trained similar embeddings on the @eveismyname and Paris 2024 datasets. We found that in most cases, notably the @eveismyname dataset, the amount of noise required for DP-SGD to achieve a reasonable value of $\varepsilon$ for DP is so high that the resulting embedding contains negligible information about the training dataset. In particular, the results for $\varepsilon=1$ are almost indistinguishable to $\varepsilon\approx 0$, as shown in Figure 5. We believe that this is simply because DP-SGD is not designed to handle such small datasets in the order of 100 images. Additional results can be found in Appendix D.
4.4.2 Differentially Private Adaptation Using Style Guidance
@@
We extend our approach to style guidance (SG) by leveraging the framework of Universal Guidance (Bansal et al., 2024). Specifically, we focus on CLIP-based style guidance, which optimizes the similarity between the CLIP embeddings of a target image and the generated image.
We encode each target image $x^{(i)}$ as $u^{(i)}$ via a CLIP image encoder, then aggregate the embeddings $u^{(1)},\dots,u^{(n)}$ into $u^{*}$ using (5) or (7), depending on whether subsampling is applied. The aggregated embedding $u^{*}$ is then incorporated into the reverse diffusion process as a style guide. Further implementation detailes are provided in Appendix C.
We apply our SG-based approach to both datasets. While it provides privacy protection by obfuscating embedding details, the resulting images captured only generalized stylistic elements and lack the detailed fidelity and coherence achieved with the TI-based method. As shown in Figure 6, this highlights the superiority of TI in balancing privacy and high-quality image generation.
The reduced effectiveness of SG for style transfer may stem from its sensitivity to hyperparameters such as the guidance weight $w$, leading to instability. Although Bansal et al. (2024) proposed remedies, namely backward guidance and per-step self-recurrence, these proved insufficient for our application. Additionally, the CLIP embeddings may not retain enough stylistic detail after the aggregation.